The owner of a movie theater was counting the money from 1 day's ticket sales. He knew that a total of 150 tickets were sold. Adult tickets cost $12 each and children's tickets cost $8 each. If the total receipts for the day was $1576. How many of each type of ticket were sold?
Answer by ikleyn(52787)

.
The owner of a movie theater was counting the money from 1 day's ticket sales. He knew that a total of 150 tickets were sold.
Adult tickets cost $12 each and children's tickets cost $8 each. If the total receipts for the day was $1576.
How many of each type of ticket were sold?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Solution 1 (one equation approach)

12*a + 8*(150-a) = 1576.

Solve for a.


Solution 2 (two equations approach)

a + c = 150,
12a + 8c = 1576.


Solve the system for "a" and "c".